Write the partial fraction decomposition of the rational expression. 11x - 48 x2 - 9x + 20
Answer #1

Solution-

We have given the rational expression as-

11.0 48 2 – 9.r + 20

Now, let us first write denominator in factor form.

\frac{11x-48}{x^2-5x-4x+20}

=\frac{11x-48}{x(x-5)-4(x-5)}

=\frac{11x-48}{(x-5)(x-4)}

Now, this can be written in the form as partial fraction as

\frac{11x-48}{(x-5)(x-4)}=\frac{A}{x-5}+\frac{B}{x-4} .....(1)

Taking lcm on right side.

\frac{11x-48}{(x-5)(x-4)}=\frac{A(x-4)+B(x-5)}{(x-5)(x-4)}

Or

Taking Denominator of the sides.

11x-48=A(x-4)+B(x-5)

On putting x =4, we get

11(4)-48=A(4-4)+B(4-5)

44-48=A(0)+B(-1)

-4=-B

4=B

Or

B=4

Again putting x = 5, we get

11(5)-48=A(5-4)+B(5-5)

55-48=A(1)+B(0)

7=A

Or

A=7

On putting A=7 and B= 4 (as calculated above) in equation (1), we get

\frac{11x-48}{(x-5)(x-4)}=\frac{7}{x-5}+\frac{4}{x-4}

Hence, the given rational expression can be written in partial fraction decomposition form as

\frac{11x-48}{x^2-9x+20}=\frac{7}{x-5}+\frac{4}{x-4}


answered by: ANURANJAN SARSAM